OpenStudy (anonymous):

Find the coordinates of all points on the graph of y =1 - r^2 at which the tangent line passes through the point (2,0)

OpenStudy (chaise):

\[y = -r^2 -1\] \[y \prime=-2r\] At r = 2 --> -4 The at (2,0) will be -4. y = mx+b 0 = -4(2)+b 0 = -8+b b = 8 y=mx+b y=-4x+8

OpenStudy (anonymous):

hmm i got it thanks :D